Reynoso Corporation manufactures titanium and aluminum tennis racquets. Reynoso’s total overhead costs consist of assembly costs and inspection costs. The following information is available:

Cost Titanium Aluminum Total Cost
Assembly 500 mach. hours 500 mach. hours $60000
Inspections 350 150 $92000
2100 labor hours 1900 labor hours



Reynoso is considering switching from one overhead rate based on labor hours to activity-based costing.
Total overhead costs assigned to titanium racquets, using a single overhead rate, are

a. $76000.

b. $92000.

c. $79800.

d. $106400.
Answer #1

Overhead rate = (60,000+92,000)/(2100+1900)

= 152,000/4000

= 38 per dlh

Overhead assigned to titanium racquets = 38*2100

= 79,800

Option C is the answer